# Synergistic eigenanalysis of covariance and Hessian matrices for enhanced binary classification on health datasets

### v2

@@ 7 Discussion
The experiments conducted on the WBCD dataset illustrate these points. As shown in Figure 9, the absolute covariance matrices for each class in the WBCD dataset reveal a pattern that approximates the isotropic condition. The relatively uniform higher values along the diagonal and generally lower values in the off-diagonal elements indicate that the covariance matrices are nearly proportional to the identity matrix. As discussed in Section 3, the WBCD dataset being normalized (as with the other datasets used in this study) helps approximate isotropy, as does the dataset being a good dataset with low redundancy among attributes. This near-isotropy supports the method’s effectiveness in maximizing the between-class mean distance when projecting the WBCD data into the leading eigenvector of the covariance matrix, as suggested by Theorem 1.
@@
Figure 9: Absolute covariance matrices for WBCD dataset by class. The figure presents the absolute values of the covariance matrices for the WBCD dataset, calculated separately for each class. These matrices highlight the strength of linear relationships between features within each class. In these absolute covariance matrices, the approximate isotropic condition is evidenced by the relatively uniform higher values along the diagonal and the generally lower values in the off-diagonal elements. Specifically, for Class 0, the average diagonal value is 0.5655, and the average non-diagonal value is 0.1566; for Class 1, these values are 0.8503 and 0.2828, respectively. These characteristics, though not starkly isotropic, approximate the conditions where the covariance matrices are proportional to the identity matrix. This approximation helps ensure the applicability of Theorem 1 in higher dimensions.
@@
In contrast, datasets that do not fit the ideal conditions of isotropy and dominant leading eigenvalues may present challenges for the proposed method. For instance, if the covariance matrix elements are random or lack a pattern approximating proportionality to the identity matrix, which can occur in unnormalized datasets or datasets with high redundancy among attributes, the method’s effectiveness may be compromised. Additionally, if the covariance and/or Hessian matrices exhibit relatively uniform eigenvalues, where the leading eigenvalues are not sufficiently dominant, the method may not perform as effectively.
@@
While the proposed method employs the LDA criteria, it surpasses LDA itself in all cases. Figures 7 and 8 visually demonstrates the advantages of the proposed method over LDA. LDA is limited to a one-dimensional projection for binary classification problems [51], where it seeks to identify a single direction that simultaneously satisfies the two criteria. Conversely, the proposed method splits the task of meeting the criteria into two directions. The utilization of higher dimensionality in the proposed method increases the likelihood of discovering class separability, aligning with Cover’s theorem [73]. KDA, operating on a non-linear mode, outperforms LDA in all cases, yet it remains fundamentally confined to one dimension, restricting its effectiveness in capturing intricate class-specific patterns compared to our proposed method.
Our method outperforms LOL by leveraging non-linear modes of operation, providing a distinct advantage in capturing complex patterns beyond the linear capabilities of LOL. UMAP exhibits good class separability with the widest margin as shown in Figures 7 and 8. However, it is important to note that UMAP is not inherently a classification technique, and thus, it fails to generalize well to new data in Figure 8 . Similar limitations apply to LLE, which, although effective in revealing local data structures, lacks the inherent capability for classification and generalization.
Figures 7 and 8 also underscores the simplicity and interpretability of linear SVMs as basic linear classifiers in dealing with low-dimensional data. The figure shows clear visualizations of the SVM’s decision boundaries and separation achieved through the various dimensionality reduction methods. Notably, our proposed method combined with linear SVMs offers valuable insights into the decision-making process within the underlying DNN. Leveraging the transparency and interpretability provided by SVMs to address the inherent opacity of DNNs, our method bridges the gap between complex DNNs and comprehensible, highly accurate decision processes within a 2D space.
@@ A.1 Proof of Theorem 1 - Maximizing Covariance for Maximizing Squared Between-Class Mean Distance
To prove that maximizing the variance will maximize the squared between-class mean distance, we start by considering two sets of 1D data points representing two classes, denoted as $C_{1}$ and $C_{2}$, each consisting of $n$ samples. The data in $C_{1}$ and $C_{2}$ follow the same underlying distribution centered around their respective means, $m_{1}$ and $m_{2}$. In other words, $C_{2}$ can be understood as a reflection of $C_{1}$ with the axis of reflection positioned at the overall mean, denoted as $m$, effectively giving rise to a shifted configuration. Specifically, the indices for $C_{1}$ range from 1 to $n$, while the indices for $C_{2}$ range from $n+1$ to $2n$.
For any pair of data points $x_{i}$ and $x_{j}$, where $1\leq i\leq n$ and $n+1\leq j\leq 2n$, the difference between $x_{i}$ and $m_{1}$ is equal to the difference between $x_{j}$ and $m_{2}$, i.e., $x_{i}-m_{1}=x_{j}-m_{2}$. This relationship ensures that the shifted copies of $C_{1}$ and $C_{2}$ maintain the same relative distances from their respective means, preserving the identical distribution in both sets. Furthermore, the variances of $C_{1}$ and $C_{2}$, are equal, denoted as $\sigma_{1}^{2}=\sigma_{2}^{2}=\sigma_{w}^{2}$, and the combined data from $C_{1}$ and $C_{2}$ has a variance of $\sigma^{2}$.
@@
The between-class mean distance, denoted as $d$, represents the separation between the means of $C_{1}$ and $C_{2}$. We can express the means as $m_{1}=m-\frac{d}{2}$ and $m_{2}=m+\frac{d}{2}$, where $m$ is effectively located in the middle, equidistant from both $m_{1}$ and $m_{2}$.
@@
The variance of the combined data is given by:
$\sigma^{2}=\frac{1}{2n-1}\left(\sum_{i=1}^{n}(x_{i}-m)^{2}+\sum_{i=n+1}^{2n}(x_{i}-m)^{2}\right)$
To simplify the expression, we substitute $m=m_{1}+\frac{1}{2}d$ and $m=m_{2}-\frac{1}{2}d$. This yields:
$s^{2}$ $=\frac{1}{2n-1}\left(\sum_{i=1}^{n}\left(x_{i}-\left(m_{1}+\frac{1}{2}d\right)\right)^{2}+\sum_{i=n+1}^{2n}\left(x_{i}-\left(m_{2}-\frac{1}{2}d\right)\right)^{2}\right)$
$=\frac{1}{2n-1}\left(\sum_{i=1}^{n}\left(\left(x_{i}-m_{1}\right)-\frac{1}{2}d\right)^{2}+\sum_{i=n+1}^{2n}\left(\left(x_{i}-m_{2}\right)+\frac{1}{2}d\right)^{2}\right)$
$\begin{split}&=\frac{1}{2n-1}\left(\sum_{i=1}^{n}\left(\left(x_{i}-m_{1}\right)^{2}-\left(x_{i}-m_{1}\right)d+\frac{1}{4}d^{2}\right)\right)\\ &+\left.\sum_{i=n+1}^{2n}\left(\left(x_{i}-m_{2}\right)^{2}+\left(x_{i}-m_{2}\right)d+\frac{1}{4}d^{2}\right)\right)\\ &=\frac{1}{2n-1}\left(\sum_{i=1}^{n}\left(\left(x_{i}-m_{1}\right)^{2}+\frac{1}{4}d^{2}\right)-\sum_{i=1}^{n}\left(\left(x_{i}-m_{1}\right)d\right)\right)\\ &+\left.\sum_{i=n+1}^{2n}\left(\left(x_{i}-m_{2}\right)^{2}+\frac{1}{4}d^{2}\right)\right)+\sum_{i=n+1}^{2n}\left(\left(x_{i}-m_{2}\right)d\right)\\ &=\frac{1}{2n-1}\left(\sum_{i=1}^{n}\left(\left(x_{i}-m_{1}\right)^{2}+\frac{1}{4}d^{2}\right)+\left.\sum_{i=n+1}^{2n}\left(\left(x_{i}-m_{2}\right)^{2}+\frac{1}{4}d^{2}\right)\right)\right)\\ &+\sum_{i=n+1}^{2n}\left(\left(x_{i}-m_{2}\right)d\right)-\sum_{i=1}^{n}\left(\left(x_{i}-m_{1}\right)d\right)\\ &=\frac{1}{2n-1}\left(\sum_{i=1}^{n}\left(\left(x_{i}-m_{1}\right)^{2}+\frac{1}{4}d^{2}\right)+\left.\sum_{i=n+1}^{2n}\left(\left(x_{i}-m_{2}\right)^{2}+\frac{1}{4}d^{2}\right)\right)\right)\\ &+\left(\sum_{i=n+1}^{2n}\left(x_{i}-m_{2}\right)-\sum_{i=1}^{n}\left(x_{i}-m_{1}\right)\right)d\\ \end{split}$
The property of the identical distribution in both sets suggests $\sum_{i=n+1}^{2n}\left(x_{i}-m_{2}\right)-\sum_{i=1}^{n}\left(x_{i}-m_{1}\right)=0$. So, we get:
$\sigma^{2}$ $=\frac{1}{2n-1}\left(\sum_{i=1}^{n}\left(\left(x_{i}-m_{1}\right)^{2}+\frac{1}{4}d^{2}\right)+\sum_{i=n+1}^{2n}\left(\left(x_{i}-m_{2}\right)^{2}+\frac{1}{4}d^{2}\right)\right)$
$=\frac{1}{2n-1}\left(\sum_{i=1}^{n}\left(x_{i}-m_{1}\right)^{2}+\sum_{i=n+1}^{2n}\left(x_{i}-m_{2}\right)^{2}+\frac{1}{2}n\cdot d^{2}\right)$
$\approx\frac{1}{2\left(n-1\right)}\left(\sum_{i=1}^{n}\left(x_{i}-m_{1}\right)^{2}+\sum_{i=n+1}^{2n}\left(x_{i}-m_{2}\right)^{2}\right)+\frac{1}{4}d^{2}$
$=\frac{1}{2}\left(\frac{\sum_{i=1}^{n}\left(x_{i}-m_{1}\right)^{2}}{\left(n-1\right)}+\frac{\sum_{i=n+1}^{2n}\left(x_{i}-m_{2}\right)^{2}}{\left(n-1\right)}\right)+\frac{1}{4}d^{2}$
$=\frac{1}{2}\left(\sigma_{w}^{2}+\sigma_{w}^{2}\right)+\frac{1}{4}d^{2}$
$=\sigma_{w}^{2}+\frac{1}{4}d^{2}.$
@@
Now, considering the data points representing the projected data onto an (Eigen)vector, we can utilize the Variance Ratio Preservation Theorem for Projection onto a Vector, which establishes the relationship between $\sigma^{2}$ and $\sigma_{w}^{2}$ as follows:
$\sigma_{w}^{2}=\lambda\cdot\sigma^{2}$
@@
where $\lambda$ is a constant between 0 and 1, determined by the distribution of the original data being projected.
@@ A.2 Proof of Theorem 2 - Maximizing Hessian for Minimizing Within-Class Variance
Hence, maximizing the Hessian matrix ($\mathrm{H}_{\theta}$) will minimize the within-class variances ($\sigma_{\text{post}}^{2}$) as desired.
A.3 Supporting theorems
A.3.1 Distance Preservation Theorem for Projection onto a Vector
In a 2-dimensional space with a first axis and a second axis, consider a set of points located solely on the first axis, denoted by $(x_{i},0)$, where $x_{i}$ represents the position of a point along the first axis. Let $\mathbf{v}=(v_{1},v_{2})$ be a unit vector, with $v_{1}$ and $v_{2}$ as its components. The projection operation maps each point $(x_{i},0)$ to its corresponding scalar projection $y_{i}$ onto the vector $\mathbf{v}$. The Distance Preservation Theorem states that the projection onto $\mathbf{v}$ preserves the ratio of distances between the points on the first axis.
Proof:
The scalar projection $y_{i}$ of a point $(x_{i},0)$ onto the vector $\mathbf{v}$ is given by:
$y_{i}=\frac{(x_{i},0)\cdot\mathbf{v}}{|\mathbf{v}|}=\frac{(x_{i},0)\cdot\mathbf{v}}{1}=x_{i}\cdot v_{1}+0\cdot v_{2}=x_{i}\cdot v_{1}$
This implies that the projection of each point $(x_{i},0)$ from the first axis onto the vector $\mathbf{v}$ is obtained by multiplying the coordinate $x_{i}$ by the first component $v_{1}$ of $\mathbf{v}$.
Now, considering the distances between two points $(x_{i},0)$ and $(x_{j},0)$ on the first axis and their corresponding projections $y_{i}$ and $y_{j}$ onto the vector $\mathbf{v}$, we define the distance between $(x_{i},0)$ and $(x_{j},0)$ as:
$d_{1}=|x_{i}-x_{j}|$
Similarly, the distance between $y_{i}$ and $y_{j}$ is defined as:
$d_{2}=|y_{i}-y_{j}|$
Substituting the expressions for $y_{i}$ and $y_{j}$ derived earlier, we have:
$d_{2}=|x_{i}\cdot v_{1}-x_{j}\cdot v_{1}|=|v_{1}|\cdot|x_{i}-x_{j}|$
It can be observed that the ratio of the distances is constant:
$\frac{d_{2}}{d_{1}}=\frac{|v_{1}|\cdot|x_{i}-x_{j}|}{|x_{i}-x_{j}|}=|v_{1}|$
This shows that the ratio of distances between the points on the first axis is preserved in the projection onto $\mathbf{v}$.
Therefore, the Distance Preservation Theorem for Projection onto a Vector concludes that the projection operation onto the vector $\mathbf{v}=(v_{1},v_{2})$ in a 2-dimensional space preserves the ratio of distances between the points on the first axis.
A.3.2 Variance Preservation Theorem for Projection onto a Vector
Consider an arbitrary subset of points located on the first axis, denoted by $X=\{x_{1},x_{2},\ldots,x_{n}\}$, with a mean $\mu_{X}$ and variance $\sigma_{X}^{2}$. These points are projected onto the vector $\mathbf{v}=(v_{1},v_{2})$ using the projection operation defined earlier.
The projected subset of points on $\mathbf{v}$ is denoted by $Y=\{y_{1},y_{2},\ldots,y_{n}\}$, where each $y_{i}$ represents the projection of $x_{i}$ onto $\mathbf{v}$. The variance of $Y$, denoted as $\sigma_{Y}^{2}$, is a measure of the spread of the projected points around its mean $\mu_{Y}$.
The Variance Preservation Theorem states that the ratio of variances between the projected subset $Y$ and the original subset $X$ is equal to the square of the first component $v_{1}$ of the projection vector $\mathbf{v}$.
Mathematically, this can be expressed as:
$\frac{\sigma_{Y}^{2}}{\sigma_{X}^{2}}=v_{1}^{2}$
Proof:
The variance of $Y$ can be computed as:
$\sigma_{Y}^{2}=\frac{1}{n-1}\sum_{i=1}^{n}(y_{i}-\mu_{Y})^{2}$
Similarly, the variance of $X$ can be computed as:
$\sigma_{X}^{2}=\frac{1}{n-1}\sum_{i=1}^{n}(x_{i}-\mu_{X})^{2}$
According to the Distance Preservation Theorem for Projection onto a Vector, the squared distances between the points on the first axis and their projections can be related as:
$(y_{i}-\mu_{Y})^{2}=|y_{i}-\mu_{Y}|^{2}=(|v_{1}|\cdot|x_{i}-\mu_{X}|)^{2}=v_{1}^{2}\cdot(x_{i}-\mu_{X})^{2}$
Substituting this expression into the variance of $Y$, we have:
$\sigma_{Y}^{2}=\frac{1}{n-1}\sum_{i=1}^{n}v_{1}^{2}\cdot(x_{i}-\mu_{X})^{2}=v_{1}^{2}\cdot\frac{1}{n-1}\sum_{i=1}^{n}(x_{i}-\mu_{X})^{2}=v_{1}^{2}\cdot\sigma_{X}^{2}$
Therefore, we have shown that:
$\sigma_{Y}^{2}=v_{1}^{2}\cdot\sigma_{X}^{2}$
Hence, the ratio of variances between the projected subset $Y$ and the original subset $X$ is given by $v_{1}^{2}$, as stated in the Variance Preservation Theorem for Projection onto a Vector.

### v3

@@ 7 Discussion
The experiments conducted on the WBCD dataset illustrate these points. As shown in Figure 8, the absolute covariance matrices for each class in the WBCD dataset reveal a pattern that approximates the isotropic condition. The relatively uniform higher values along the diagonal and generally lower values in the off-diagonal elements indicate that the covariance matrices are nearly proportional to the identity matrix. In contrast, as shown in D, the absolute within-class covariance matrices for the unnormalized WBCD dataset exhibit a significant dominance of a few components that overshadow the other components, resulting in a structure that is far from isotropic and a far less suitable foundation for achieving effective class separability. This highlights the importance of normalization as discussed in Subsection 3.2. The WBCD dataset being normalized (as with the other datasets used in this study) helps approximate isotropy, as does the dataset being a well-structured dataset with low redundancy among attributes. This near-isotropy supports the method’s effectiveness in maximizing the between-class mean distance when projecting the multidimensional data into the leading eigenvector of the covariance matrix, as suggested by Theorem 1. Notably, the superior performance of the proposed method under these near-isotropic conditions demonstrates its robustness and ability to achieve optimal results even when perfect isotropy is not strictly present.
@@
The dominance of the leading eigenvectors in capturing the most significant information is further reflected in the comparative performance results shown in Figure 10. The results show that while Hesssian method and the proposed method sometimes outperform the standard DNNs, in other cases, DNNs perform better. This variability arises because DNNs utilize the entire feature space, while the projection methods operate within a reduced eigenspace using only a subset of the most informative eigendirections. When the omitted directions contain meaningful information, the standard DNNs excel; conversely, when the excluded features are primarily noise, the projection methods yield superior results. Importantly, in all scenarios, the proposed method consistently achieves strong performance, highlighting its ability to leverage the dominant eigenvectors and capture significant information from both the covariance and Hessian matrices.
Figure 8: Absolute covariance matrices for (normalized) WBCD dataset by class. The figure presents the absolute values of the covariance matrices for the WBCD dataset, calculated separately for each class. These matrices highlight the strength of linear relationships between features within each class. In these absolute covariance matrices, the approximate isotropic condition is evidenced by the relatively uniform higher values along the diagonal and the generally lower values in the off-diagonal elements. Specifically, for Class 0, the average diagonal value is 0.5655, and the average non-diagonal value is 0.1566; for Class 1, these values are 0.8503 and 0.2828, respectively. These characteristics, though not starkly isotropic, approximate the conditions where the covariance matrices are proportional to the identity matrix. This approximation helps ensure the applicability of Theorem 1 in higher dimensions, as evidenced by the superior performance of the proposed method in practice.
@@
Figure 10: Performance comparison of DNN and the two DNN-based data projection methods using cross-validation on four distinct datasets: (a) WBCD, (b) heart disease, (c) neural spike train, and (d) Pima Indians diabetes datasets. This figure presents the average F1 score, ROC AUC, and Cohen’s Kappa values obtained through 5- or 10-fold cross-validation for DNNs and the two DNN-based projection techniques: Hessian method and the proposed method. Both methods are referred to as DNN-based methods because they rely on the eigendirections derived from the Hessian matrix evaluated on a trained deep neural network. In subfigures (b) and (c), the DNN-based projection methods outperform the standard DNNs, while in subfigures (a) and (d), the standard DNNs show superior performance. This difference arises because DNNs utilize the entire feature space, while the projection methods operate within a reduced eigenspace, leveraging only one or two key eigendirections. When the excluded eigendirections in the full space contain additional relevant information, the DNNs outperform the projection methods; conversely, when the excluded features are primarily noise, the projection methods yield better performance by focusing on the most informative directions.
In contrast, datasets that do not fit the ideal conditions of isotropy and dominant leading eigenvalues may present challenges for the proposed method. For instance, if the within-class covariance matrix elements are random or lack a pattern approximating proportionality to the identity matrix, which can occur in unnormalized datasets or datasets with high redundancy among attributes, the method’s effectiveness may be compromised. Additionally, if the covariance and/or Hessian matrices exhibit relatively uniform eigenvalues, where the leading eigenvalues are not sufficiently dominant, the method may not perform as effectively.
@@
While the proposed method employs the LDA criteria, it surpasses LDA itself in all cases. Figures 6 and 7 visually demonstrates the advantages of the proposed method over LDA. LDA is limited to a one-dimensional projection for binary classification problems [51], where it seeks to identify a single direction that simultaneously satisfies the two criteria. Conversely, the proposed method splits the task of meeting the criteria into two directions. The utilization of higher dimensionality in the proposed method increases the likelihood of discovering class separability, aligning with Cover’s theorem [80]. KDA, operating on a non-linear mode, outperforms LDA in all cases, yet it remains fundamentally confined to one dimension, restricting its effectiveness in capturing intricate class-specific patterns compared to our proposed method.
Our method outperforms LOL by leveraging non-linear modes of operation, providing a distinct advantage in capturing complex patterns beyond the linear capabilities of LOL. UMAP exhibits good class separability with the widest margin as shown in Figures 6 and 7. However, it is important to note that UMAP is not inherently a classification technique, and thus, it fails to generalize well to new data in Figure 7 . Similar limitations apply to LLE, which, although effective in revealing local data structures, lacks the inherent capability for classification and generalization.
Figures 6 and 7 also underscores the simplicity and interpretability of linear SVMs as basic linear classifiers in dealing with low-dimensional data. The figure shows clear visualizations of the SVM’s decision boundaries and separation achieved through the various dimensionality reduction methods. Notably, our proposed method combined with linear SVMs offers valuable insights into the decision-making process within the underlying DNN. It projects data into a space defined by the leading Hessian direction, which captures the key curvature directions in the DNN’s loss landscape. This allows us to visualize which features have the greatest impact on the DNN’s decision-making process (Figure 5, right subfigure). SVM decision boundaries then illustrate class separation in this reduced space (Figures 6 and 7, bottom right subfigures). This combined approach bridges the gap between complex decision boundaries of DNNs and comprehensible, highly accurate decision processes within a 2D space.
@@ A.1 Proof of Theorem 1 - Maximizing covariance for maximizing squared between-class mean distance
To prove that maximizing the variance will maximize the squared between-class mean distance, we start by considering two sets of 1D data points representing two classes, denoted as $C_{1}$ and $C_{2}$, each consisting of $n$ samples. The data in $C_{1}$ and $C_{2}$ are centered around their respective means, $\mu_{1}$ and $\mu_{2}$. Let $\mu$ denote the overall mean of the combined data, expressed as:
$\mu=\frac{1}{2n}\left(\sum_{i=1}^{n}x_{i}+\sum_{i=n+1}^{2n}x_{i}\right).$
@@
The between-class mean distance, denoted as $d$, represents the separation between the means of $C_{1}$ and $C_{2}$. We can express the means as $\mu_{1}=\mu-\frac{d}{2}$ and $\mu_{2}=\mu+\frac{d}{2}$, where $\mu$ is effectively located in the middle, equidistant from both $\mu_{1}$ and $\mu_{2}$.
@@
Furthermore, the variances of $C_{1}$ and $C_{2}$, are equal, denoted as $\sigma_{1}^{2}=\sigma_{2}^{2}=\sigma_{w}^{2}$, and the combined data from $C_{1}$ and $C_{2}$ has a variance of $\sigma^{2}$. The variance of the combined data is given by:
$\sigma^{2}=\frac{1}{2n-1}\left(\sum_{i=1}^{n}(x_{i}-\mu)^{2}+\sum_{i=n+1}^{2n}(x_{i}-\mu)^{2}\right)$
To simplify the expression, we substitute $\mu=\mu_{1}+\frac{1}{2}d$ and $\mu=\mu_{2}-\frac{1}{2}d$. This yields:
$s^{2}$ $=\frac{1}{2n-1}\left(\sum_{i=1}^{n}\left(x_{i}-\left(\mu_{1}+\frac{1}{2}d\right)\right)^{2}+\sum_{i=n+1}^{2n}\left(x_{i}-\left(\mu_{2}-\frac{1}{2}d\right)\right)^{2}\right)$
$=\frac{1}{2n-1}\left(\sum_{i=1}^{n}\left(\left(x_{i}-\mu_{1}\right)-\frac{1}{2}d\right)^{2}+\sum_{i=n+1}^{2n}\left(\left(x_{i}-\mu_{2}\right)+\frac{1}{2}d\right)^{2}\right)$
$\begin{split}&=\frac{1}{2n-1}\left(\sum_{i=1}^{n}\left(\left(x_{i}-\mu_{1}\right)^{2}-\left(x_{i}-\mu_{1}\right)d+\frac{1}{4}d^{2}\right)\right)\\ &+\left.\sum_{i=n+1}^{2n}\left(\left(x_{i}-\mu_{2}\right)^{2}+\left(x_{i}-\mu_{2}\right)d+\frac{1}{4}d^{2}\right)\right)\\ &=\frac{1}{2n-1}\left(\sum_{i=1}^{n}\left(\left(x_{i}-\mu_{1}\right)^{2}+\frac{1}{4}d^{2}\right)-\sum_{i=1}^{n}\left(\left(x_{i}-\mu_{1}\right)d\right)\right)\\ &+\left.\sum_{i=n+1}^{2n}\left(\left(x_{i}-\mu_{2}\right)^{2}+\frac{1}{4}d^{2}\right)\right)+\sum_{i=n+1}^{2n}\left(\left(x_{i}-\mu_{2}\right)d\right)\\ &=\frac{1}{2n-1}\left(\sum_{i=1}^{n}\left(\left(x_{i}-\mu_{1}\right)^{2}+\frac{1}{4}d^{2}\right)+\left.\sum_{i=n+1}^{2n}\left(\left(x_{i}-\mu_{2}\right)^{2}+\frac{1}{4}d^{2}\right)\right)\right)\\ &+\sum_{i=n+1}^{2n}\left(\left(x_{i}-\mu_{2}\right)d\right)-\sum_{i=1}^{n}\left(\left(x_{i}-\mu_{1}\right)d\right)\\ &=\frac{1}{2n-1}\left(\sum_{i=1}^{n}\left(\left(x_{i}-\mu_{1}\right)^{2}+\frac{1}{4}d^{2}\right)+\left.\sum_{i=n+1}^{2n}\left(\left(x_{i}-\mu_{2}\right)^{2}+\frac{1}{4}d^{2}\right)\right)\right)\\ &+\left(\sum_{i=n+1}^{2n}\left(x_{i}-\mu_{2}\right)-\sum_{i=1}^{n}\left(x_{i}-\mu_{1}\right)\right)d\\ \end{split}$
Now, let’s consider the properties of the two subsets $C_{1}$ and $C_{2}$. By the zero-sum property, we have:
$\sum_{i=1}^{n}(x_{i}-\mu_{1})=0\quad\text{and}\quad\sum_{i=n+1}^{2n}(x_{i}-\mu_{2})=0.$
Substitute these properties into the variance equation:
$\sigma^{2}$ $=\frac{1}{2n-1}\left(\sum_{i=1}^{n}\left(\left(x_{i}-\mu_{1}\right)^{2}+\frac{1}{4}d^{2}\right)+\sum_{i=n+1}^{2n}\left(\left(x_{i}-\mu_{2}\right)^{2}+\frac{1}{4}d^{2}\right)\right)$
$=\frac{1}{2n-1}\left(\sum_{i=1}^{n}\left(x_{i}-\mu_{1}\right)^{2}+\sum_{i=n+1}^{2n}\left(x_{i}-\mu_{2}\right)^{2}+\frac{1}{2}n\cdot d^{2}\right)$
$\approx\frac{1}{2\left(n-1\right)}\left(\sum_{i=1}^{n}\left(x_{i}-\mu_{1}\right)^{2}+\sum_{i=n+1}^{2n}\left(x_{i}-\mu_{2}\right)^{2}\right)+\frac{1}{4}d^{2}$
$=\frac{1}{2}\left(\frac{\sum_{i=1}^{n}\left(x_{i}-\mu_{1}\right)^{2}}{\left(n-1\right)}+\frac{\sum_{i=n+1}^{2n}\left(x_{i}-\mu_{2}\right)^{2}}{\left(n-1\right)}\right)+\frac{1}{4}d^{2}$
$=\frac{1}{2}\left(\sigma_{1}^{2}+\sigma_{2}^{2}\right)+\frac{1}{4}d^{2}.$
@@
Now, considering the data points representing the projected data onto an (Eigen)vector, we can utilize the variance ratio preservation property for projection onto a vector, which establishes the relationship between $\sigma^{2}$, $\sigma_{1}^{2}$, and $\sigma_{2}^{2}$ as follows:
$\lambda_{1}=\frac{\sigma_{1}^{2}}{\sigma^{2}},\quad\lambda_{2}=\frac{\sigma_{2}^{2}}{\sigma^{2}},$
or equivalently,
$\sigma_{1}^{2}=\lambda_{1}\cdot\sigma^{2},\quad\sigma_{2}^{2}=\lambda_{2}\cdot\sigma^{2},$
@@
where each of $\lambda_{1}$ and $\lambda_{2}$ is a constant between 0 and 1, determined by the distribution of the original data being projected.
@@
$\sigma^{2}$ $=\frac{1}{2}\left(\lambda_{1}+\lambda_{2}\right)\sigma^{2}+\frac{1}{4}d^{2}.$
Let $\lambda=\frac{1}{2}\left(\lambda_{1}+\lambda_{2}\right)$. Then:
@@ A.2 Proof of Theorem 2 - Maximizing hessian for minimizing within-class variance
Hence, maximizing the Hessian ($\mathrm{H}_{\theta}$) will minimize the within-class variances ($\sigma_{\text{post}}^{2}$) as desired.
A.3 Variance ratio preservation property for projection onto a vector
The property is formalized in Theorem 4, which states that the ratio of variances between any two subsets of 1D data is preserved when the data is projected onto a 2-dimensional vector.